OpenStudy (anonymous):

@Chemistry 4C3H5(NO3)3-->12CO2+10H20+6N2+O2 How many grams of carbon dioxide are produced by the explosion of 21.0 grams of nitroglcerine?

OpenStudy (anonymous):

n = m/M. n = mols of the substance. m = mass of the substance. M = molar mass of the substance. Now what you want to do is find the mols of nitroglycerine. So, n = m/M. Now we need m and M. m = 21.0 g (in the question). M = ? To find M, look at the periodic table or use your memory to determine it. For example N corresponds to 14.0 g/mol. After you have found n, look at the equation. You see that for every single of nitroglycerine you have 12 CO2 particles. So find n for nitroglycerine, and multiply that by 12. Rearranging the equation: n = m/M We want m of CO2. m = n x M. Find M of CO2 (44.0 g/mol). Do the calculation.